Common Management Systems

CI Records Upgrade Completed

This page contains the current project status of the CI Records Upgrade Project currently under way.  Please see below for most recent project update postings.  This site will be updated every 4-6 hours during the upgrade period (Thursday 10/13 - Friday 10/21) until the upgrade has been completed.

Click on the links below to go directly to a specific date for status or scroll through the entries.

 Next scheduled update - 10/18/2011 at  12:00 Noon.

Tuesday 10/18/2011

Update 12:00 Noon

  • Setup of the new nightly batch schedule is currently in progress and will be completed and scheduled for tonight, 10/18.
  • Data validation for ALL functional areas has been completed and they have been cleared for turnover.
  • 100% of ASG configuration tasks have been completed.
  • 100% of ASG migration tasks have been completed.
  • 100% of ASG Security tasks have been completed.
  • CI Records is now available via myCI.

Monday 10/17/2011

Update 04:00 PM 

  • Setup of the new nightly batch schedule is currently in progress and will be completed and scheduled for tomorrow night, 10/18.
  • Data validation for Faculty Affair data will be completed tomorrow morning, 10/18.
  • Data validation for all other functional areas has been completed.
  • 100% of ASG configuration tasks have been completed.
  • 100% of ASG migration tasks have been completed.
  • 100% of ASG Security tasks have been completed.
  • CI Records rollout is being scheduled for 10/18 at Noon pending final data validation from Faculty Affairs. 

Update 10:00 AM 

  • Setup of the new nightly batch schedule is currently in progress.
  • Setup and configuration for Academic Advising has been completed.
  • Student Finance has completed their data validation for Student Financials.
  • Budget completed their data validation for LCD.
  • Admissions and Records completed their data validation for the Admissions, Records, and CS portion of Campus Community.
  • Setup and configuration for Admissions and Student Records is completed.
  • Data validation for all functional areas is currently in progress.
  • 100% of ASG configuration tasks have been completed.
  • 100% of ASG migration tasks have been completed.
  • 100% of ASG Security tasks have been completed.
  • The HR issue related to position data has been resolved..
  • We remain on schedule

Update 12:00 AM

  • 100% of ASG configuration tasks have been completed.
  • 100% of ASG migration tasks have been completed.
  • 100% of ASG Security tasks have been completed.
  • Setup of the new nightly batch schedule is scheduled for tomorrow, 10/17.
  • Student Finance has completed their data validation for Student Financials.
  • Budget completed their data validation for LCD.
  • Admissions and Records completed their data validation for the Admissions, Records, and CS portion of Campus Community.
  • Setup and configuration for Admissions and Student Records is completed.
  • Setup and configuration for Academic Advising is scheduled for tomorrow, 10/17.
  • Data validation for all functional areas is scheduled for tomorrow, 10/17.
  • We are currently working on an HR issue related to position data with the CO.
  • We remain on schedule

Sunday 10/16/2011

Update 08:00 PM

  • 100% of ASG configuration tasks have been completed.
  • 97% of ASG migration tasks have been completed.
  • 64% of ASG Security tasks have been completed.
  • ASG continues applying new security and validating migrations.
  • Setup of the new nightly batch schedule is scheduled for tomorrow, 10/17.
  • Admissions and Records completed their data validation for the Admissions, Records, and CS portion of Campus Community.
  • Setup and configuration for Admissions and Student Records is completed.
  • Setup and configuration for Academic Advising is scheduled for tomorrow, 10/17.
  • Data validation for all other functional areas is scheduled for tomorrow, 10/17.
  • We remain on schedule.

Update 04:00 PM

  • 100% of ASG configuration tasks have been completed.
  • 66% of ASG migration tasks have been completed.
  • 42% of ASG Security tasks have been completed.
  • ASG continues their portion of the move to production tasks which are estimated to be completed some time this evening. 
  • System setup and configuration has started by Admissions and Records.
  • A couple of issues had come up but were quickly resolved with assistance from the CO.
  • We remain on schedule.

Update 10:00 AM

  • 77% of ASG configuration tasks have been completed.
  • 18% of ASG migration tasks have been completed
  • ASG continues their portion of the move to production tasks which are estimated to be completed some time this evening. 

Saturday 10/15/2011

Update 06:30 PM

  • ASG continues their portion of the move to production tasks which are estimated to be completed some time tomorrow evening. 

Update 01:30 PM

  • The CO has completed their portion of the upgrade process and has turned the production instance over to ASG.
  • ASG has begun their portion of the move to production tasks which are estimated to be completed some time tomorrow evening. 
  • We are currently ahead of schedule.

Update 08:00 AM

  • The CO has completed the primary upgrade steps and is running our final back up.
  • The CO should be ready to turn the production instance back to ASG in about 6 hours.
  • We are currently ahead of schedule

Friday 10/14/2011

Update 11:00 PM:

  • The CO continues working on the application data conversions and updates. 

Update 05:00 PM:

  • ASG has completed required data extracts from HCIPRJ.
  • The CO has successfully completed work on the PeopleTools upgrade.
  • The CO is now working on the application data conversions and updates. 

Update 01:00 PM:

  • The CO completed the restore of the initial backup into HCIPRJ at 10:00AM, ahead of schedule.
  • The CO continues work on the PeopleTools upgrade. 
  • ASG has started data extracts from HCIPRJ.

Update 08:00 AM:

  • The CO continues the restore of the initial backup into HCIPRJ.
  • The CO continues work on the PeopleTools upgrade. 

Thursday 10/13/2011

Update 11:30 PM:

  • The Chancellor's Office (CO) has completed initial database environmental tasks.
  • The CO is working on the PeopleTools upgrade now. 
  • Unisys is working on the restore of the initial backup into HCIPRJ. Estimated completion of this task is Noon tomorrow, 10/14.

Update 06:00 PM: 

  • The cloning process was started for non-production instance HCITRS at 01:00PM.
  • Final production batch processing was completed as scheduled.
  • Functional pre-upgrade tasks were started and completed as scheduled.
  • The cloning process for the non-production instance HCITRS was completed ahead of schedule.  Security was updated, and the instance has been turned over to our Program Mangers.  
  • The CashNet Pre-Upgrade activities have been completed.
  • The production instance of CI Records (HCIPRD) was turned over to the Chancellor's Office for upgrade at 06:00PM.

Update 12:00 PM: 

  • Cloning has completed for the non-production instance HCITST and the instance has been turned over to Academic Advising.  
  • The next cloning process, for non-production instance HCITRS, is scheduled to begin at 01:00PM.  Estimated completion of this task is 05:00 PM today.
  • The myCI link to CI Records has been turned off and the redirect message has been turned on.  This will be in place for the duration of the upgrade process.

Update 07:00 AM: 

  • Cloning process of current production instance (HCIPRD) to non-production instance HCITST has started.  Estimated completion of this task is 12:00 PM today.

Back to CMS Page

 

©